# Responsible AI Question Bank: A Comprehensive Tool for AI Risk Assessment

### v1

@@ 5.2 RAI deep dive assessment for investors
Case study overview. We utilized the RAI Question Bank for the integration of Environment, Social and Governance (ESG) and AI for investors. This project aimed to develop the ESG-AI investor framework 1, which was implemented from April 2023 to April 2024. Collaborative research [26] was applied for this project, allowing researchers and practitioners to work together in developing the framework. The practitioners involved in this project included not only investors as potential users, but also senior design experts and industrial experts in both AI and ESG. The following outlines the key roles and responsibilities defined in the project.
  • AI researcher: Provide insights into AI principles, technologies, and responsible AI practices; Drive the framework and toolkit development.
  • ESG expert/investor: Share perspectives on investment strategies, ESG considerations and assessment needs.
  • Design expert: Ensure the framework’s usability from a user-centered design perspective.
  • Industrial expert: Contribute industry-specific knowledge and practical insights while balancing out different perspective from AI researchers and investors.
  • Project manager: Coordinate the project activities and kept the research team on schedule.
The main creators of the RAI Question Bank were involved in this project as AI researchers and played a crucial role in developing the ESG-AI framework by applying the RAI Question Bank.
@@
AI use case. As we live in the Industry 4.0 era, AI is considered a crucial technology for enhancing performance, product quality, and employee well-being [27]. This shift has led to the widespread use of AI applications across all sectors of daily life and industry. To conduct preliminary screening and identify companies with significant AI use cases, investors must understand how AI is being utilized within the sectors in their investment portfolios. The AI use case component is designed to address this need. This component can be used to identify high-impact AI use cases across various industry sectors from the investor’s perspective.
RAI governance indicators. This component was developed using insights from various sources, including interviews with 28 global and regional companies, public company disclosures (e.g., annual reports), and the needs identified by investors during several workshops conducted for this project. It consists of 10 high-level indicators designed to assess a company’s overall commitment, accountability, and measurement of RAI. These indicators are categorized into four areas: Board Oversight, RAI Commitment, RAI Implementation, and RAI Metrics.
RAI deep dive assessment. This component is developed to facilitate detailed analysis and engagement with company management on AI governance and RAI practices. AI ethics principles encompass key values and guidelines that address RAI development and deployment. Investors can use this assessment for a systematic evaluation of fairness, transparency, accountability, privacy, and more, contributing to a holistic understanding of how well a company adheres to ethical standards in its AI practices.
@@
Through collaboration with practitioners, we reviewed and tailored these questions through multiple iterations to suit the project’s specific needs. The final version of the RAI deep dive assessment includes 42 sub-questions across the eight AI ethics principles. Additionally, we incorporated RAI metrics to support investors’ assessments and to encourage company transparency and public disclosure. As discussed earlier, these metrics are specifically included to inject quality aspects into the questions and provide evidence beyond minimal yes/no answers.
Table 3 shows the example questions and metrics selected for Accountability.
@@
User feedback and impact on the community. During the project, we conducted multiple workshops with investors (i.e., potential users) and senior industrial experts, engaging in iterative testing and improvements. These dedicated participants contributed to the ongoing review process, providing valuable feedback on the framework’s usability, particularly the risk assessment questions from the user perspective.
The user feedback focused mainly on the understandability (readability) and usability of the questions and metrics in real-world scenarios and the clarity of the guidance provided for each assessment item. Feedback included recommendations such as:
  • Restructuring the question bank and adding a question for each AI ethics principle:
@@
    This change aimed to provide high-level insights and a better understanding of RAI practices and the risks associated with each principle. The previous version of the RAI Question Bank did not include principle-specific questions; however, based on this feedback, we modified the structure. The current version now includes principle questions and corresponding sub-questions.
  • Improving the clarity of questions and metrics that might be difficult for general users to understand:
@@
    In response, we identified and revised questions containing technical or specialized terms to enhance their quality and ensure they are more accessible. We also revised the metrics, including their descriptions, to provide clearer guidance and improve the overall assessment process. These changes aim to ensure that the questions and metrics are not only comprehensible to a wider audience but also maintain a high standard of quality and relevance.
  • Instructions for users on how to evaluate the answers of examinees:
@@
    This aims to enhance risk assessors’ capability in accurately judging responses, ensuring consistent and effective assessments. These guidelines can help assessors better understand the criteria for evaluation, leading to more reliable and insightful risk assessments. Additionally, they assist users in addressing the quality aspects of questions by incorporating relevant metrics, beyond relying on simple yes/no answers. We have added specialized instructions for users into the framework, providing guidance on evaluating the quality of evidence and the context of responses. However, due to the large scale of the RAI Question Bank, this approach has not yet been applied to the entire question set. We recognize the importance of these enhancements and will consider and discuss further improvements for future iterations.
This iterative process ensured that the risk questions in the framework was refined to meet the needs of its users effectively. As a result, the RAI Question Bank-based questions was recognized as a highly useful and practical tool for AI and ESG risk assessment for investors.
@@
The framework was publicly released in April 2023. Within the first 24 hours, it received coverage from 23 media outlets, with numbers continuing to rise. In the first week, over 1,000 people downloaded our final report from the website, and around 100 downloaded the framework toolkit, despite the gated website requiring personal information. The framework also gained significant attention on social media, where it was praised for providing valuable insights into addressing key challenges faced by practitioners and for going beyond a principles-only approach to implementing RAI.
@@ 6 Discussion
In this section, we demonstrate practical uses of the RAI Question Bank, which may be of interest to practitioners aiming to develop and use AI ethically. Specifically, we demonstrate how the RAI Question Bank can help companies examine their RAI practices to comply with key AI regulations and standards such as the EU AI Act. We also discuss the application of the RAI Question Bank for AI agents as foundation model-based AI agents have been widely adopted in this area.
@@ 6.1 Compliance with regulations
In recent years, many countries have been working to develop AI bills and regulations to ensure the secure use of AI. For example, the EU AI Act has become the first comprehensive regulation on AI in March 2024. In the US, AI bill of rights  [^23] https://www.whitehouse.gov/ostp/ai-bill-of-rights/ developed by the White House Office of Science and Technology Policy (OSTP) in 2022, , serves as a blueprint rather than a set of laws and is therefore regarded as a voluntary regulation. Bill C-27  [^24] https://www.parl.ca/legisinfo/en/bill/44-1/c-27 in Canada, also known as the Digital Charter Implementation Act, is a proposed law currently making its way through the Canadian Parliament. Introduced in April 2022, it is currently undergoing review. In South Korea, Bill on AI Liability was suggested in March 2023, but like other regulations, it has not yet been passed into law. The Australian government has been drafting the AI Safety Standard in collaboration with many experts in the field, which has not been publicly released yet.
These regulations commonly include specific requirements and considerations for high-risk AI. Consequently, there are growing concerns among companies developing and using AI about compliance with current and upcoming regulations, especially for high-risk AI. To support this, we illustrate how practitioners can utilize the RAI Question Bank to examine their RAI practices against the legal requirements identified in the EU AI Act, based on our mapping approach depicted below.
High-risk AI refers to AI systems that may pose significant risks to the health and safety or fundamental rights of natural persons. Examples include AI systems used for biometrics, critical infrastructure and access to essential services. Well-known use cases of high-risk AI include credit scoring systems and facial recognition systems.
We present the following three-steps approach to assess high-risk AI systems against the EU AI Act.
  • Step1. Identify legal requirements from the EU AI Act:
    Begin by thoroughly reviewing the EU AI Act to identify all relevant legal requirements for high-risk AI systems. This step involves understanding the specific obligations and standards set by the regulation.
  • Step2. Map the requirements with risk questions in the RAI Question Bank:
    Next, take the identified legal requirements and map them to the risk questions provided in the RAI Question Bank. This mapping process ensures that each legal requirement is addressed by corresponding risk assessments, creating a comprehensive compliance checklist.
  • Step3. Score the AI system:
    Finally, evaluate the AI system by scoring it against the mapped risk questions. This scoring process helps determine the extent to which the AI system meets the legal requirements, highlighting areas of compliance and identifying gaps that need to be addressed.
Identify legal requirements from the EU AI Act. We conducted a thorough review of the EU AI Act to identify the specific legal requirements that apply to high-risk AI systems. This involved analyzing the text of the Act to extract all relevant legal requirements that developers of high-risk AI systems must comply with. Through this process, we identified 21 key requirements that address various aspects of AI ethics, including transparency, accountability, fairness, and privacy.
@@
These requirements are designed to ensure that high-risk AI systems do not pose significant risks to the health, safety, or fundamental rights of individuals. They include provisions for proper documentation, risk management, human oversight, data governance, and more. This step is crucial for understanding the regulatory landscape and establishing a foundation for systematic risk assessment and management of high-risk AI systems.
Table 4 presents 10 example requirements that we identified for high-risk AI systems.
Category ID Requirement AI principle
@@
Risk management E01 Risk management system shall be established, implemented, documented and maintained. P8
@@
Data governance and management E02 For datasets, should consider relevant data preparation, prior quality assessment, examination (e.g., biases), possible data gaps or shortcoming. P4
@@
Technical specification E03 General description (e.g., intended purpose) and detailed description (e.g., design specification, architecture). P6
Record keeping E04 Record keeping of all relevant documentation and information for traceability. P8
@@
Transparency E05 User instructions including provider details, capabilities and limitations, purposes, performance, specifications, etc. P6
@@
Transparency E06 If the AI system is interacting with humans, the users should be informed that they are interacting with an AI system and the content has been artificially generated or manipulated. P6
@@
Human oversight E07 Monitor its operation, consider over-reliance, override or reverse the output and intervene on the operation or interupt the system (e.g., through a “stop” button). P2
Cybersecurity E08 Cybersecurity and resilience to prevent and control for attacks. P4
Register E09 Register their systems in an EU-wide database in any external AI database (e.g., EU-wide database). P6
Compliance E10 Conformity assessment and management process. P8
AI principle:
P1: Human, Societal, Environmental Wellbeing, P2: Human-centred Values, P3: Fairness, P4: Privacy and security
P5: Reliability and safety, P6: Transparency and explainabilty, P7: Contestability, P8: Accountability
@@
Table 4: 10 examples requirements identified from the EU AI Act with a short summary; the details can be found in the EU AI Act document.
@@
Map the requirements with risk questions in the RAI Question Bank. By mapping these requirements to specific risk questions in the RAI Question Bank, we aim to provide a comprehensive framework for assessing compliance and ensuring that AI systems are developed and used responsibly. This step involves a detailed analysis of each identified requirement from the EU AI Act and its alignment with the ethical principles and risk factors outlined in the RAI Question Bank. Accordingly, each requirement was carefully interpreted to understand its implications and scope. We then matched each requirement with one or more relevant questions from the RAI Question Bank that address similar concerns or objectives. This mapping process ensures that for every legal requirement, there is a corresponding risk question that can be used to evaluate whether an AI system meets the necessary requirements.
For example, a requirement related to technical specification might be mapped to questions about the AI system’s transparency and explainability, the documentation of its decision-making processes, and the accessibility of its outputs to users. Similarly, a requirement focused on data management would be mapped to questions about data privacy, security measures, and data management practices.
@@
For this mapping, we designed a simple schema as shown in Figure 13. In the figure, we also show example screenshots for each entity.
Figure 13: The designed schema for mapping the requirements to the RAI Question Bank questions.
  • EU-base is a core entity that includes the identifier (ID), requirement description and corresponding section in the document.
  • EU-QB represents a mapping table that comprises the requirement ID and question ID. Each requirement has at least one or multiple mappings in this table.
  • EU-QB-sub includes two question IDs to demonstrate subsequent questions. Each question in EU-QB can have zero or one mapping.
This schema is straightforward yet facilitates effective navigation of risk questions that cover all compliance requirements.
Score the AI system. Users can provide answers to the questions using a simple binary metric based on their situation.
@@
For the question, Do you establish an AI risk management system to conduct ongoing risk assessment and treatment?, users can answer Yes if they have implemented a structured system for ongoing assessment and treatment of AI risks. Answering No indicates the absence of such a system. N/A should be selected if the organization does not engage in AI activities or if this specific practice is not applicable to their operations. Scores are assigned based on whether the answer reflects compliance (Yes) or non-compliance (No), with N/A indicating exemption from scoring.
This approach provides clarity on how to interpret and score responses in a compliance assessment using a simple binary metric. Adjustments can be made based on specific scoring criteria or compliance frameworks relevant to the user context.
@@ 6.2 Empowering trust of AI agents
This section addresses another potential and promising area, AI agents where the RAI Question Bank can play a focal role in the AI space.
As introduced earlier, AI agents can autonomously complete user tasks by interacting with users and other parties, such as external foundation models, tools, and other agents [28]. Unlike typical AI systems, these agents share more control with humans (e.g., the agent owner, users) and can operate autonomously without direct human intervention. For example, an agent can use a credit card to make purchases on behalf of the user. This example underscores the need for responsible development and use of AI agents, with ongoing risk assessments to prevent serious incidents and harms to individuals, society, and the environment [7]. Additionally, when users set specific goals for the agent, the underlying foundation model is queried to achieve these goals. It is crucial to prevent the foundation model from being influenced by adversarial inputs and to ensure it does not generate harmful or undesirable outputs for users or other components [15]. To foster trust in AI agents, both the agents themselves and the foundation models they interact with must be trustworthy. the RAI Question Bank can be applied to the both sides. Figure 14 shows the conceptual diagram for this application.
Figure 14: The potential application of the RAI Question Bank for foundation model-based AI agents.
Risk assessment of AI agents. We have adopted RAI plugins from the AI agent design patterns proposed in 2023 [7]. RAI plugins compriese six agent components: Continuous risk assessor, Black box recorder, Explainer, Multimodal guardrails, and AIBOM registry. These components serve as an RAI role to assess and monitor risks and to record all relevant data and information of the AI agent. We have interpreted the roles and requirements of these components and identified relevant risk questions from the RAI Question Bank as follows (Table 5).
@@ 13. Does the AI agent verify the provenance of the components and manage their execution? -
Source:
ISO- ISO/IEC AI management
EU- EU Assessment List for Trustworthy Artificial Intelligence
AIA- Canada Algorithmic Impact Assessment
NIST- NIST AI Risk Management Framework 1.0
@@
The Black Box Recorder is an essential component for ensuring the traceability of AI systems. It serves the crucial function of recording and storing comprehensive information about the AI system’s operations, decisions, and performance. This includes logging inputs, outputs, processing steps, and any anomalies encountered during the system’s operation. By maintaining a detailed record, it supports accountability and facilitates audits, making it possible to review and understand the system’s behavior retrospectively. More importantly, this component enhances transparency for stakeholders, enabling them to verify that the AI system operates as intended and adheres to relevant regulatory and ethical standards. Additionally, it helps in assigning clear responsibility by documenting who interacted with the system and how, thereby contributing to a robust governance framework. There are three questions for this component adapted from EU Assessment List for Trustworthy Artificial Intelligence and Canada Algorithmic Impact Assessment framework.
@@
Explainer plays a crucial role in providing important information about the AI agent, including its roles, capabilities, limitations, and outputs. This component is specifically designed to enhance transparency and explainability practices. Key considerations for this component include that AI systems should produce understandable outcomes, that explainability should be integrated from the very early design stages, and that all necessary information should be thoroughly documented. By addressing these points, the Explainer ensures that stakeholders have a clear understanding of how the AI system operates and can trust its decisions.
@@
Multimodal guardrail is designed to prevent inappropriate multimodal inputs from being sent to the foundation model and to monitor and control the outputs from the model. It should be capable of processing various forms of data, such as text, audio, and video, to support comprehensive control [15]. Positioned in the middle layer between the foundation model and other components, this component serves a crucial role in the operational loop.
@@
AIBOM Registry refers to the storage for all necessary information associated with the AI supply chain, including procurement information. The AIBOM registry can include not only internal components and their information but also external tools, agents, and models. To ensure RAI practices, the AI agent needs to verify the provenance of the components. Accordingly, this component is strongly interrelated with the Black Box Recorder.
@@
Risk assessment of foundation models. As a key external entity, a foundation model generates outputs in response to the AI agent’s prompts. To mitigate risks and minimize negative impacts on AI agents, model providers should conduct ongoing risk assessments and ensure compliance with regulations and standards.
From the AI agent’s perspective, it is essential to have confidence that the foundation model is developed and deployed responsibly. To support this, we have identified a set of essential requirements for verifying foundation models, primarily derived from the EU AI Act. Additionally, we suggest relevant questions based on the RAI Question Bank to check these requirements.
Table 6 shows the requirement list and corresponding questions to address the requirements from the AI agent’s perspective.

### v2

@@ 5.2 RAI Deep Dive Assessment for Investors
Case study overview. We utilized the RAI Question Bank for the integration of Environment, Social and Governance (ESG) and AI for investors. This project aimed to develop the ESG-AI investor framework [2], which was implemented from April 2023 to April 2024. Collaborative research [87] was applied for this project, allowing researchers and practitioners to work together in developing the framework. The practitioners involved in this project included not only investors as potential users, but also senior design experts and industrial experts in both AI and ESG.
@@
Through collaboration with ESG experts, we reviewed and tailored these questions through multiple iterations to suit the project’s specific needs. The final version of the RAI deep dive assessment includes 42 sub-questions across the eight AI ethics principles. Additionally, we incorporated RAI metrics to support investors’ assessment and encourage company transparency and public disclosure. As discussed earlier, these metrics are specifically included to inject quality aspects into the questions and provide evidence beyond minimal yes/no answers. Table 2 shows the example questions and metrics selected for Accountability.
@@
User feedback and impact on the community. During the project, we conducted multiple workshops with investors (i.e., potential users) and senior industrial experts, engaging in iterative testing and improvements. These dedicated participants contributed to the ongoing review process, providing valuable feedback on the framework’s usability, particularly the risk assessment questions from the user perspective. Accordingly, user feedback primarily addressed the readability and usability of the questions and metrics in real-world scenarios, as well as the clarity of the guidance for each assessment item. Feedback included recommendations such as:
@@
Restructuring the question bank and adding a question for each AI ethics principle: This change aimed to provide high-level insights and a better understanding of RAI practices and the risks associated with each principle. The previous version of the RAI Question Bank did not include principle-specific questions; however, based on this feedback, we modified the structure. The current version now includes principle questions and corresponding sub-questions.
@@
Improving the clarity of questions and metrics that might be difficult for general users to understand: In response, we identified and revised questions containing technical or specialized terms to enhance their quality and ensure they are more accessible. We also revised the metrics, including their descriptions, to provide clearer guidance and improve the overall assessment process. These changes aim to ensure that the questions and metrics are not only comprehensible to a wider audience but also maintain a high standard of quality and relevance.
@@
Instructions for users on how to evaluate the answers of examinees: This aims to enhance risk assessors’ capability in accurately judging responses, ensuring consistent and effective assessments. These guidelines can help assessors better understand the criteria for evaluation, leading to more reliable and insightful risk assessments. They assist users in addressing the quality aspects of questions by incorporating relevant metrics, beyond relying on simple yes/no answers. We have added specialized instructions for users into the framework, providing guidance on evaluating the quality of evidence and the context of responses. Due to the large scale of the RAI Question Bank, this approach has not yet been applied to the entire question set. We recognize the importance of these enhancements and will consider and discuss further improvements for future iterations.
@@
This iterative process ensured that the risk questions in the framework was refined to meet the needs of its users effectively. As a result, the RAI Question Bank-based questions was recognized as a highly useful and practical tool for AI and ESG risk assessment for investors. The framework was publicly released in April 2023. Within the first 24 hours, it received coverage from 23 media outlets. In the first week, over 1,000 people downloaded our final report from the website, and around 100 downloaded the framework toolkit, despite the gated website requiring personal information. The framework also gained significant attention on social media, where it was praised for providing valuable insights into addressing key challenges faced by practitioners and for going beyond a principles-only approach to implementing RAI.
@@ 6 Discussion
In this section, we demonstrate practical uses of the RAI Question Bank, which may be of interest to practitioners aiming to develop and use AI ethically. Specifically, we demonstrate how the RAI Question Bank can help companies examine their RAI practices to comply with key AI regulations and standards such as the EU AI Act. We also discuss the application of the RAI Question Bank for foundation model-based AI agents, given their widespread adoption in this area [88].
@@ 6.1 Compliance with Regulations
In recent years, many countries have been working to develop AI bills and regulations to ensure the secure use of AI. For example, the EU AI Act has become the first comprehensive regulation on AI in March 2024. In the US, AI bill of rights [89] developed by the White House Office of Science and Technology Policy (OSTP) in 2022, serves as a blueprint regarded as a voluntary regulation. Bill C-27 [90] in Canada, also known as the Digital Charter Implementation Act, is a proposed law currently making its way through the Canadian Parliament. In South Korea, Bill on AI Liability was suggested in March 2023. It has not yet been passed into law. The Australian government has released the Voluntary AI Safety Standard [91] and National framework for the assurance of AI in government [92] in 2024. These regulations commonly include specific requirements for high-risk AI as it may pose significant risks to the health and safety or fundamental rights of natural persons (e.g., credit scoring [93] and facial recognition [94]). Companies developing and using AI are concerning about compliance with current and upcoming regulations [95]. To support this, we illustrate how practitioners can utilize the RAI Question Bank to examine their RAI practices against the legal requirements identified in the EU AI Act, based on our mapping approach depicted below.
@@
1) Identify legal requirements from the EU AI Act. We conducted a thorough review of the EU AI Act to identify the specific legal requirements that apply to high-risk AI systems. This involved analyzing the text of the Act to extract all relevant legal requirements that developers of high-risk AI systems must comply with. Through this process, we identified 21 key requirements that address various aspects of AI ethics, including transparency, accountability, fairness, and privacy. These requirements are designed to ensure that high-risk AI systems do not pose significant risks to the health, safety, or fundamental rights of individuals. They include provisions for proper documentation, risk management, human oversight, data governance, and more. This step is crucial for understanding the regulatory landscape and establishing a foundation for systematic risk assessment and management of high-risk AI systems. Table 3 presents 10 example requirements that we identified for high-risk AI systems.
Category ID Requirement
@@
Risk management E01 Risk management system shall be established, implemented, documented and maintained.
@@
Data governance and management E02 For datasets, should consider relevant data preparation, prior quality assessment, examination (e.g., biases), possible data gaps or shortcoming.
@@
Technical specification E03 General description (e.g., intended purpose) and detailed description (e.g., design specification, architecture).
Record keeping E04 Record keeping of all relevant documentation and information for traceability.
@@
Transparency E05 User instructions including provider details, capabilities and limitations, purposes, performance, specifications, etc.
@@
Transparency E06 If the AI system is interacting with humans, the users should be informed that they are interacting with an AI system and the content has been artificially generated or manipulated.
@@
Human oversight E07 Monitor its operation, consider over-reliance, override or reverse the output and intervene on the operation or interupt the system (e.g., through a “stop” button).
Cybersecurity E08 Cybersecurity and resilience to prevent and control for attacks.
Register E09 Register their systems in an EU-wide database in any external AI database (e.g., EU-wide database).
Compliance E10 Conformity assessment and management process.
@@
Table 3: 10 examples requirements identified from the EU AI Act with a short summary; the details can be found in the EU AI Act document [21].
@@
2) Map the requirements with risk questions in the RAI Question Bank. By mapping these requirements to specific questions in the RAI Question Bank, we aim to provide a comprehensive framework for assessing compliance and ensuring that AI systems are developed and used responsibly. This step involves a detailed analysis of each requirement from the Act and its alignment with the ethical principles and risk factors outlined in the RAI Question Bank. Accordingly, each requirement was carefully interpreted to understand its implications and scope. We then matched each requirement with one or more relevant questions from the RAI Question Bank that address similar concerns or objectives. This mapping process ensures that for every legal requirement, there is a corresponding risk question that can be used to evaluate whether an AI system meets the necessary requirements. For example, a requirement related to technical specification might be mapped to questions about the AI system’s transparency and explainability, the documentation of its decision-making processes, and the accessibility of its outputs to users. Similarly, a requirement focused on data management would be mapped to questions about data privacy, security measures, and data management practices.
@@
3) Score the AI system. Users can provide answers to the questions using a simple binary metric based on their situation.
@@
For the question, Do you establish an AI risk management system to conduct ongoing risk assessment and treatment?, users can answer Yes if they have implemented a structured system for ongoing assessment and treatment of AI risks. Answering No indicates the absence of such a system. N/A should be selected if the organization does not engage in AI activities or if this specific practice is not applicable to their operations. Scores are assigned based on whether the answer reflects compliance (Yes) or non-compliance (No), with N/A indicating exemption from scoring. This approach provides clarity on how to interpret and score responses in a compliance assessment using a simple binary metric. Adjustments can be made based on specific scoring criteria or compliance frameworks relevant to the user context.
@@ 6.2 Empowering Trust of AI Agents
AI agents can autonomously complete user tasks by interacting with users and other parties, such as external foundation models, tools, and other agents [96]. These agents share more control with humans (e.g., the agent owner, users) and can operate autonomously without direct human intervention (e.g., use of a credit card to make purchases on behalf of the user). This underscores the need for responsible development and use of AI agents, with ongoing risk assessments to prevent serious incidents and harms to individuals, society, and the environment [97]. It is crucial to prevent the foundation model from being influenced by adversarial inputs and ensure it does not generate harmful or undesirable outputs for users or other components [29]. To foster trust in AI agents, both the agents themselves and the foundation models they interact with must be trustworthy. The RAI Question Bank can be applied to the both sides.
Risk assessment of AI agents. We have adopted RAI plugins from the AI agent design patterns proposed in 2023 [97]. RAI plugins comprises five agent components: Continuous risk assessor, Black box recorder, Explainer, Multimodal guardrails, and AIBOM registry. These components serve as an RAI role to assess and monitor risks and to record all relevant data and information of the AI agent. We have interpreted the roles and requirements of these components and identified relevant risk questions from the RAI Question Bank as follows (Table 4).
@@ 13. Does the AI agent verify the provenance of the components and manage their execution? -
The Black Box Recorder is an essential component for ensuring the traceability of AI systems [98]. It serves the crucial function of recording and storing comprehensive information about the AI system’s operations, decisions, and performance. This includes logging inputs, outputs, processing steps, and any anomalies encountered during the system’s operation. By maintaining a detailed record, it supports accountability and facilitates audits, making it possible to review and understand the system’s behavior retrospectively. More importantly, this component enhances transparency for stakeholders, enabling them to verify that the AI system operates as intended and adheres to relevant regulatory and ethical standards [99]. Additionally, it helps in assigning clear responsibility by documenting who interacted with the system and how, thereby contributing to a robust governance framework [100]. There are three questions for this component adapted from EU Assessment List for Trustworthy Artificial Intelligence and Canada Algorithmic Impact Assessment framework.
@@
Explainer plays a crucial role in providing important information about the AI agent, including its roles, capabilities, limitations, and outputs. This component is specifically designed to enhance transparency and explainability practices. Key considerations for this component include that AI systems should produce understandable outcomes, that explainability should be integrated from the very early design stages, and that all necessary information should be thoroughly documented [101]. By addressing these points, the Explainer ensures that stakeholders have a clear understanding of how the AI system operates and can trust its decisions.
@@
Multimodal guardrail is designed to prevent inappropriate multimodal inputs from being sent to the foundation model and to monitor and control the outputs from the model [102]. It should be capable of processing various forms of data, such as text, audio, and video, to support comprehensive control [29]. Positioned in the middle layer between the foundation model and other components, this component serves a crucial role in the operational loop.
@@
AIBOM Registry refers to the storage for all necessary information associated with the AI supply chain, including procurement information [103]. The AIBOM registry can include not only internal components and their information but also external tools, agents, and models [104]. To ensure RAI practices, the AI agent needs to verify the provenance of the components. Accordingly, this component is strongly interrelated with the Black Box Recorder.
@@
Risk assessment of foundation models. As a key external entity, a foundation model generates outputs in response to the AI agent’s prompts. To mitigate risks and minimize negative impacts on AI agents, model providers should conduct ongoing risk assessments and ensure compliance with regulations and standards [105]. From the AI agent’s perspective, it is essential to have confidence that the foundation model is developed and deployed responsibly. To support this, we have identified a set of essential requirements for verifying foundation models, primarily derived from the EU AI Act. Table 5 shows the requirement list and corresponding questions to address the requirements from the AI agent’s perspective.